# MAT-DiSMech: A Discrete Differential Geometry-based Computational Tool for Simulation of Rods, Shells, and Soft Robots

### v1

@@
    Accurate and efficient simulation tools are essential in robotics, enabling the visualization of system dynamics and the validation of control laws before committing resources to physical experimentation. Developing physically accurate simulation tools is particularly challenging in soft robotics, largely due to the prevalence of geometrically nonlinear deformation. A variety of robot simulators tackle this challenge by using simplified modeling techniques – such as lumped mass models – which lead to physical inaccuracies in real-world applications. On the other hand, high-fidelity simulation methods for soft structures, like finite element analysis, offer increased accuracy but lead to higher computational costs. In light of this, we present a Discrete Differential Geometry-based simulator that provides a balance between physical accuracy and computational speed. Building on an extensive body of research on rod and shell-based representations of soft robots, our tool provides a pathway to accurately model soft robots in a computationally tractable manner. Our open-source MATLAB-based framework is capable of simulating the deformations of rods, shells, and their combinations, primarily utilizing implicit integration techniques. The software design is modular for the user to customize the code – for example, add new external forces and impose boundary conditions– to suit their requirements. The implementations for prevalent forces encountered in robotics, including gravity, contact, kinetic and viscous friction, and aerodynamic drag, have been provided. We provide several illustrative examples that showcase the capabilities and validate the physical accuracy of the simulator. The open-source code is available at https://github.com/StructuresComp/dismech-matlab. We anticipate that the proposed simulator can serve as an effective digital twin tool, enhancing the Sim2Real pathway in soft robotics research.
@@ 1.  Introduction
However, accurately modeling the dynamic behavior of soft robots – that, in principle, have infinite degrees of freedom – presents significant challenges [4]. Unlike traditional rigid robots, which follow well-understood principles of kinematics and rigid body mechanics, soft robots rely on materials governed by the principles of continuum mechanics, enabling them to undergo deformations that are inherently nonlinear. These nonlinearities arise from both geometry and material properties. Geometric nonlinearity results from deformations in slender structures such as bending, twisting, or stretching, where the relationship between displacements and strains becomes nonlinear. Material nonlinearity, by contrast, is attributed to nonlinear relationship between stress and strain. In this paper, we focus on addressing geometric nonlinearity.
@@
Figure 1: Examples of robots or soft structures: Pneunet actuator, Manta ray, and Parachute, modeled using MAT-DiSMech. The key features within MAT-DiSMech. (a) Elasticity: The framework accounts for stretching, bending and twisting deformation modes for elastic rods and stretching, bending deformations for shells. (b) External forces: Some external forces implemented in MAT-DiSMech are shown. Additionally, the codebase allows users to implement their own external forces. (c) Self contact: Self contact is modeled using IMC [20, 35] which treats contact implicitly by defining a penalty energy as a function of $\Delta$, the minimum distance between two edges that can potentially come in contact. (d) Actuation: Actuation is primarily done by modifying the natural curvature or twist of the elastic elements. The parameters, $\bar{\kappa},\bar{\tau}$, and $\bar{\phi}$, denote the natural curvature of the rod, the natural twist of the rod, and the natural hinge angle of the shell, respectively.
@@ 2.  Dynamical Modeling
Degrees of Freedom and Kinematics. To model the dynamics of an elastic rod or a network of rods, we use the DER method [18], which builds on the Kirchhoff rod theory [36]. According to this method, the rod can undergo three kinds of deformations: stretching, bending, and twisting. The degrees of freedom (DOF) vector for rods comprises the coordinates of the nodes, denoted by $\mathbf{x}$, and the twist angles of the edges, denoted by $\theta$, to keep track of their rotation. Unlike rods, shells can undergo only two types of deformation: stretching or membrane deformation and bending. We implement two distinct discrete elastic shell theories to model shell bending, allowing the user to select their preferred method: (1) Hinge-based [23] and (2) Mid-edge normal-based [24]. For hinge-based bending, the DOF vector for shell only consists of the nodal positions. Therefore, for a structure with $N$ nodes and $E$ edges, the DOF vector is of size $3N+E$ and is expressed as
@@
where superscript represents the transpose operator. When using the mid-edge normal-based theory, it is necessary to track the edges on the shells—similar to the edges on the rods—by introducing an additional scalar degree of freedom, denoted by $\xi$, for each of the shell-edges. The definition of $\xi$ is explained in Section 3.3.2. Given a shell with $T\times 3$ triangle array, the shell-edges can be obtained from the triangle array. Hence, if using the mid-edge normal-based formulation, assuming $Z$ to be the number of shell-edges obtained as above, the DOF vector is of size $3N+E+Z$ and is expressed as
@@
where $\mathbf{M}$ is a square diagonal lumped mass matrix of size $3N+E+Z$, dot $\dot{(\,)}$ represents derivative with respect to time, the vector $\mathbf{F}^{\textrm{elastic}}=-\frac{\partial E^{\textrm{elastic}}}{\partial\mathbf{q}}$ represents the elastic force (where $E^{\textrm{elastic}}$ is the elastic energy), $\mathbf{F}^{\text{ext}}$ is the external force vector (such as gravity), and $\mathbf{F}^{\textrm{IMC}}$ represents the contact and friction forces following the implicit contact method (IMC) [35, 20].
@@
$\mathbf{u}_{k+1}=\frac{\mathbf{q}_{k+1}-\mathbf{q}_{k}}{\Delta t}$ (5)
@@ 3.2.  Bending and Twisting Energy
The scalar curvatures along the first and second material directors using the curvature binomial are
@@
An edge may be shared by multiple bending-twisting springs, especially while modeling a network of rods. The above formulation assumes a specific orientation for the edges ($\mathbf{x}_{m}\to\mathbf{x}_{n}$ for the first edge and $\mathbf{x}_{n}\to\mathbf{x}_{o}$ for the second edge). As shown in Figure 2(g), for bending-twisting spring $\{\mathbf{x}_{a},\mathbf{x}_{b},\mathbf{x}_{c}\text{ and }\mathbf{e}^{i},\mathbf{e}^{j}\}$, since both rod-edges $\mathbf{e}^{i}$ and $\mathbf{e}^{j}$ are pointing towards the node $\mathbf{x}_{b}$, the negative of $\mathbf{e}^{j}$ is used for calculations so that it points away from and $\mathbf{e}^{i}$ points towards the node $\mathbf{x}_{b}$. The reference frame vector $\mathbf{d}^{j}_{1}$, material frame vector $\mathbf{m}^{j}_{1}$, and edge rotation angle $\theta^{j}$ are multiplied by $-1$ for computing the bending and twisting forces. Additionally, the resulting force vector and the Jacobian are reoriented to their original states by multiplying the affected terms by $-1$ if the edge and its related quantities were adjusted during the computation. Note that we could have equivalently flipped the edge $\mathbf{e}^{i}$ instead of $\mathbf{e}^{j}$ and proceeded as above, and the result would be mathematically equivalent. In the simulator, bending-twisting springs are initialized during the problem setup, and information about whether an edge needs to be negated is stored at the beginning of the simulation. This straightforward bookkeeping approach enables efficient handling of rod-like network structures.
3.3.  Shell Bending Energy
@@ 3.3.2  Mid-edge Normal-based Bending Energy
where $k\in\{p,q,r\}$ denotes the $k$-th edge of the triangle, $\mathbf{n}_{i}$ denotes the unit normal to the triangle, $\bar{A}_{i}$ denotes the undeformed area of the triangle, $\|\mathbf{\bar{e}}^{k}\|$ denotes the undeformed length of the $k$-th edge, $\mathbf{t}^{k}$ is the tangent to the surface perpendicular to the edge (note that the $\hat{(\,)}$ operator indicates unit vector), $s^{k}$ is a variable that takes values $1$ or $-1$ depending on the ownership of the mid-edge normal, and the operator $\otimes$ denotes the outer product. Figure 2(e) shows all the above-defined vectors. Note that the “shape operator” defined in Eq. (19) is not operating on a point. Instead, it refers to the $(3\times 3)$ matrix that represents the linear map from the shape operator definition in Eq. (18). For the $i$-th triangle, the discrete shell bending energy, in terms of the discrete shape operator $\Lambda_{i}$, is
@@ 5.2.  Contact and Friction with Ground
Coulomb friction force due to the ground is computed using the above contact force similar to friction force due to self contact (see Eq. (23)), except the velocity ($\hat{\mathbf{u}}$) in this case, is the velocity of the node in contact with the ground along the tangent to the ground. By default, ground is assumed to be parallel to the $x-y$ plane; hence $\hat{\mathbf{n}}$ is $[0,0,1]^{\top}$. Note that the above penalty energy based method for ground contact force is especially useful to model contact when friction is non-zero. For frictionless contact, a simpler predictor-corrector approach can be taken similar to Ref. [40] that ensures no penetration into the ground.
5.3.  Viscous Friction
@@
5.4.  Aerodynamic Drag
@@
Note that $\mathbf{J}^{\text{drag}}_{ij}$ will be non-zero only for $j=i$ and the indices $j$ that correspond to the nodes that share a triangle with the $i$-th node. The expressions for $\frac{\partial\hat{\mathbf{n}}_{k}}{\partial\mathbf{x}_{j}}$ and $\frac{\partial\mathbf{u}_{i}}{\partial\mathbf{x}_{j}}$ are provided in the Appendix B.
@@ 6.2.  Time Stepping
We employ backward Euler method for implicit time integration which allows larger time steps in comparison to explicit time integration methods. However, using backward euler leads to an artificial numerical damping of vibrations in the structure. If the vibrations are of significance in a particular application, the integration scheme can be easily replaced to suit the user requirements. While performing soft robot structures in contact-rich scenarios, the error for implicit solving may fail or take a large no. of iterations to converge. To aid in convergence in scenarios like this, we employ a line search method to get a step size which ensures that the error decreases in every iteration.
@@ 7.1.  Simulation Experiments
Figure 4: PneuNet actuator modeled as a discrete elastic rod. (a) Snapshots of the PneuNet actuator with natural curvature $15.7$ m-1 with only gravity (top) and with gravity and point force on the free end (bottom). (b) Snapshots of the PneuNet actuator with natural curvature $31.45$ m-1 with only gravity (top) and with gravity and point force on the free end (bottom). (c) Snapshots of the PneuNet actuator with natural curvature $47.15$ m-1 with only gravity (top) and with gravity and point force on the free end (bottom). Note that the cross-sectional radius of the rod is $1$ mm, it has been enlarged in the figure for better visualization. (d) The plot of the natural shapes of the rods in (a), (b), and (c) under gravity. (e) Plot of the deformed shape of the rods under gravity and point force at the free end.
@@
For our third showcase, we simulate a parachute which is modeled as a combination of elastic rods and shell. The canopy is modeled as an initially flat plate in the shape of a regular hexagon of side length $1$ m and thickness $1$ mm. The ropes are modeled as elastic rods. A rigid payload mass of $0.13$ kg is attached at the end of the ropes, which is modeled simply by increasing the mass at the node shared between all rods. The material parameters used in this simulation are: $E_{\textrm{rod}}=10$ MPa, $E_{\textrm{shell}}=1$ GPa, $\nu_{\textrm{rod}}=0.5,\nu_{\textrm{shell}}=0.3$ and $\rho=1500$ kg/m3 for both rod and shell. The external forces, in this case, are gravity and aerodynamic drag. We use $\mathbf{g}=[0,0,-9.8]^{\top}$ m/s2, and, $\rho_{\textrm{med}}=1$ kg/m3 and $C_{D}=10$ to calculate the gravity and drag force respectively. Figure 6 shows the snapshots of the parachute as it falls to the ground. It can be seen that as the simulation progresses, the initially flat canopy takes a 3D inverted bowl-like shape as expected.
@@
For our final example, we simulate a rod falling onto the ground from a height of $5$ cm. The density of the rod’s material and its Poisson ratio are taken to be $1500$ kg/m3 and $0.3$, respectively. We simulate the rod for two different values of Young’s modulus $Y$: (a) $Y=2$ MPa and (b) $Y=2$ GPa. The external forces, in this case, are gravity, ground contact, and friction. The values of $[0,0,-9.8]^{\top}$ m/s2, $0.25$, and $20$ are used respectively for the acceleration due to gravity $\mathbf{g}$, and the friction coefficient $\mu$ and contact stiffness $k_{c}$ between the ground and the rod. Figure 7 shows snapshots of the rod as it hits the ground. As Young’s modulus for the second case is much higher than the first case, we observe significant deformation of the rod in the first case while almost no deformation in the second.
Figure 7: Elastic rod with different Young’s Moduli dropped on the ground. (a1-f1) Snapshots of the elastic rod with Young’s modulus $2$ MPa, and (a2-f2) Snapshots of the elastic rod with Young’s modulus $2$ GPa. Here time is given in seconds. Note that the cross-sectional radius of the rod is $1$ mm, it has been enlarged in the figure for better visualization. (g) The trajectory of the $z$-coordinate of the lowermost point on the rod with time. The time stamps corresponding to the snapshots (a-f) are marked in the plot using vertical dashed lines.

### v2

@@
    Accurate and efficient simulation tools are essential in robotics, enabling the visualization of system dynamics and the validation of control laws before committing resources to physical experimentation. Developing physically accurate simulation tools is particularly challenging in soft robotics, largely due to the prevalence of geometrically nonlinear deformation. A variety of robot simulators tackle this challenge by using simplified modeling techniques – such as lumped mass models – which lead to physical inaccuracies in real-world applications. On the other hand, high-fidelity simulation methods for soft structures, like finite element analysis, offer increased accuracy but lead to higher computational costs. In light of this, we present a Discrete Differential Geometry-based simulator that provides a balance between physical accuracy and computational efficiency. Building on an extensive body of research on rod and shell-based representations of soft robots, our tool provides a pathway to accurately model soft robots in a computationally tractable manner. Our open-source MATLAB-based framework is capable of simulating the deformations of rods, shells, and their combinations, primarily utilizing implicit integration techniques. The software design is modular for the user to customize the code – for example, add new external forces and impose boundary conditions– to suit their requirements. The implementations for prevalent forces encountered in robotics, including gravity, contact, kinetic and viscous damping, and hydrodynamic and aerodynamic drag, have been provided. We provide several illustrative examples that showcase the capabilities and validate the physical accuracy of the simulator. The open-source code is available at https://github.com/StructuresComp/dismech-matlab. We anticipate that the proposed simulator can serve as an effective digital twin tool, enhancing the Sim2Real pathway in soft robotics research.
@@ 1.  Introduction
However, accurately modeling the dynamics of soft robots – which, in principle, possess infinite degrees of freedom – presents significant challenges [4] due to their highly nonlinear behavior. Unlike rigid body dynamics, where nonlinearity arises primarily from rotational kinematics while the body itself remains undeformed, soft body dynamics exhibit nonlinearities stemming from both rotations and large deformations of the continuum. These nonlinearities originate from two sources: geometry and material properties. Geometric nonlinearity arises when slender structures undergo significant bending, twisting, or stretching, leading to a nonlinear relationship between displacements and strains. In contrast, material nonlinearity refers to the nonlinear relationship between stress and strain inherent to the material’s constitutive behavior. In this paper, we focus on addressing geometric nonlinearity.
@@
Figure 1: Examples of robots or soft structures: Snake, Manta ray, and Parachute, modeled using MAT-DiSMech. The key features within MAT-DiSMech. (a) Elasticity: The framework accounts for stretching, bending and twisting deformation modes for elastic rods and stretching, bending deformations for shells. (b) External forces: Some external forces implemented in MAT-DiSMech are shown. Additionally, the codebase allows users to implement their own external forces. (c) Self contact: Self contact is modeled using IMC [20, 35] which treats contact implicitly by defining a penalty energy as a function of $\Delta$, the minimum distance between two edges that can potentially come in contact. (d) Actuation: Actuation is primarily done by modifying the natural curvature or twist of the elastic elements. The parameters, $\bar{\kappa},\bar{\tau}$, and $\bar{\phi}$, denote the natural curvature of the rod, the natural twist of the rod, and the natural hinge angle of the shell, respectively.
@@ 2.  Dynamical Modeling
Degrees of Freedom and Kinematics. To model the dynamics of an elastic rod or a network of rods, we use the DER method [18], which builds upon the Kirchhoff rod theory [36]. In this framework, rods can undergo three primary deformation modes: stretching, bending, and twisting. While the original Kirchhoff rod model describes inextensible rods, the DER method extends this by introducing a stretching energy, allowing the rods to deform axially. If strict inextensibility is required, it is also possible to impose edge-length-preserving constraints, for instance, using the fast projection method [37].
The degrees of freedom (DOF) vector for rods comprises the nodal positions, denoted by $\mathbf{x}$, and the twist angles of the edges, denoted by $\theta$, which capture their rotational state. Unlike rods, shells can undergo only two types of deformation: stretching or membrane deformation and bending. We implement two distinct discrete elastic shell theories to model shell bending, allowing the user to select their preferred method: (1) Hinge-based [23] and (2) Mid-edge normal-based [24]. For hinge-based bending, the DOF vector for the shell only consists of the nodal positions. Therefore, for a structure with $N$ nodes and $E$ edges, the DOF vector is of size $3N+E$ and is expressed as
@@
where superscript represents the transpose operator. When using the mid-edge normal-based theory, it is necessary to track the edges on the shells – similar to the edges on the rods – by introducing an additional scalar degree of freedom, denoted by $\xi$, for each of the shell-edges. The definition of $\xi$ is explained in Section 3.3.2. Given a shell with $T\times 3$ triangle array, the shell-edges can be obtained from the triangle array. Hence, if using the mid-edge normal-based formulation, assuming $Z$ to be the number of shell-edges obtained as above, the DOF vector is of size $3N+E+Z$ and is expressed as
@@
[…]
@@
$\mathbf{u}_{k+1}=\frac{\mathbf{q}_{k+1}-\mathbf{q}_{k}}{\Delta t}$ . (5)
@@ 3.2.  Bending and Twisting Energy
The scalar curvatures along the first and second material directors using the curvature binormal are
@@
An edge may be shared by multiple bending-twisting springs, especially while modeling a network of rods. The above formulation assumes a specific orientation for the edges ($\mathbf{x}_{m}\to\mathbf{x}_{n}$ for the first edge and $\mathbf{x}_{n}\to\mathbf{x}_{o}$ for the second edge). As shown in Figure 2(g), for bending-twisting spring $\{\mathbf{x}_{a},\mathbf{x}_{b},\mathbf{x}_{c}\text{ and }\mathbf{e}^{i},\mathbf{e}^{j}\}$, since both rod-edges $\mathbf{e}^{i}$ and $\mathbf{e}^{j}$ are pointing towards the node $\mathbf{x}_{b}$, the negative of $\mathbf{e}^{j}$ is used for calculations so that it points away from and $\mathbf{e}^{i}$ points towards the node $\mathbf{x}_{b}$. The reference frame vector $\mathbf{d}^{j}_{1}$, material frame vector $\mathbf{m}^{j}_{1}$, and edge rotation angle $\theta^{j}$ are multiplied by 1-1 for computing the bending and twisting forces. Additionally, the resulting force vector and the Jacobian are reoriented to their original states by multiplying the affected terms by 1-1 if the edge and its related quantities were adjusted during the computation. Note that we could have equivalently flipped the edge $\mathbf{e}^{i}$ instead of $\mathbf{e}^{j}$ and proceeded as above, and the result would be mathematically equivalent. In the simulator, bending-twisting springs are initialized during the problem setup, and information about whether an edge needs to be negated is stored at the beginning of the simulation. This straightforward bookkeeping approach enables efficient handling of rod-like network structures.
$3.$  Shell Bending Energy
@@ 3.3.2  Mid-edge Normal-based Bending Energy
where $k\in\{p,q,r\}$ denotes the $k$-th edge of the triangle, $\mathbf{n}_{i}$ denotes the unit normal to the triangle, $\bar{A}_{i}$ denotes the undeformed area of the triangle, $\|\mathbf{\bar{e}}^{k}\|$ denotes the undeformed length of the $k$-th edge, $\mathbf{t}^{k}$ is the tangent to the surface perpendicular to the edge (note that the $\hat{(\,)}$ operator indicates unit vector), $s^{k}$ is a variable that takes values $1$ or 1-1 depending on the ownership of the mid-edge normal, and the operator $\otimes$ denotes the outer product. Figure 2(e) shows all the above-defined vectors. Note that the “shape operator” defined in Eq. (19) is not operating on a point. Instead, it refers to the $(3\times 3)$ matrix that represents the linear map from the shape operator definition in Eq. (18). For the $i$-th triangle, the discrete shell bending energy, in terms of the discrete shape operator $\Lambda_{i}$, is
@@ 5.2.  Contact and Friction with Ground
Coulomb friction force due to the ground is computed using the above contact force similar to friction force due to self contact (see Eq. (23)), except the velocity ($\hat{\mathbf{u}}$) in this case, is the velocity of the node in contact with the ground along the tangent to the ground. By default, ground is assumed to be parallel to the xyx-y plane; hence $\hat{\mathbf{n}}$ is $0,0,1]^{\top}$. Note that the above penalty energy based method for ground contact force is especially useful to model contact when friction is non-zero. For frictionless contact, a simpler predictor-corrector approach can be taken similar to Ref. [41] that ensures no penetration into the ground.
5.3.  Viscous Damping
@@
5.4.  Hydrodynamic force using Resistive Force Theory
Resistive Force Theory (RFT) [42] is among the most widely used hydrodynamic models for describing interactions between low-Reynolds-number flows and slender structures. For the $i$-th node, the hydrodynamic force at $n$-th time instant is calculated by summing contributions from all edges connected to it, as follows,
$\mathbf{F}^{\text{RFT}}_{i}=\sum_{k\in E}-\frac{l^{k}}{2}\left[(C_{t}-C_{n})\hat{\mathbf{e}}^{k}(\hat{\mathbf{e}}^{k})^{\top}+C_{n}\mathbb{I}\right]\mathbf{u}_{i},$ (29)
[…]
$\begin{split}\mathbf{J}^{\text{RFT}}_{ij}&=\frac{\partial\mathbf{F}^{\textrm{RFT}}_{i}}{\partial\mathbf{x}_{j}}\\ &=\sum_{k\in E}-\frac{l^{k}}{2}\left[(C_{t}-C_{n})(\mathbf{u}_{i}^{\top}{\hat{\mathbf{e}}^{k}})\frac{\partial\hat{\mathbf{e}}^{k}}{\partial x_{j}}\right.\\ &\quad+(C_{t}-C_{n})\hat{\mathbf{e}}^{k}\left((\hat{\mathbf{e}}^{k})^{\top}\frac{\partial\mathbf{u}_{i}}{\partial x_{j}}+\mathbf{u}_{i}^{\top}\frac{\partial\hat{\mathbf{e}}^{k}}{\partial x_{j}}\right)\\ &\quad+\left.C_{n}\frac{\partial\mathbf{u}_{i}}{\partial x_{j}}\right],\end{split}$ (30)
where $\mathbb{I}$ is the identity matrix of size $(3\times 3)$. Note that $\mathbf{J}^{\text{RFT}}_{ij}$ will be non-zero only for $j=i$ and the indices $j$ that correspond to the nodes adjacent to the $i$-th node. The expressions for $\frac{\partial\hat{\mathbf{e}}^{k}}{\partial\mathbf{x}_{j}}$ and $\frac{\partial\mathbf{u}_{i}}{\partial\mathbf{x}_{j}}$ are provided in the Appendix B.
$5.$  Aerodynamic Drag
@@ 5.4.  Hydrodynamic force using Resistive Force Theory
Note that $\mathbf{J}^{\text{drag}}_{ij}$ will be non-zero only for $j=i$ and the indices $j$ that correspond to the nodes that share a triangle with the $i$-th node. The expressions for $\frac{\partial\hat{\mathbf{n}}_{k}}{\partial\mathbf{x}_{j}}$ and $\frac{\partial\mathbf{u}_{i}}{\partial\mathbf{x}_{j}}$ are provided in the Appendix C.
@@ 6.2.  Time Stepping
We employ the backward Euler method, an implicit time integration scheme that enables the use of larger time steps compared to explicit methods. However, backward Euler introduces significant artificial numerical damping, which can overly suppress vibrations in the structure. To address this, we also offer an implicit midpoint method, which helps reduce numerical damping and is preferable when accurate vibration behavior is important for a specific application.
Implicit integration provides substantial computational advantages over explicit schemes, particularly for systems with higher Young’s moduli. For context, Choi et al. [34] present a comparison of computational times between simulators employing explicit and implicit time-stepping methods. Nevertheless, for very simple systems with low Young’s moduli, explicit methods such as forward Euler can be faster, as they avoid the computational overhead of computing and inverting Jacobians. Therefore, we also include an implementation of the forward Euler method for users who prefer an explicit scheme.
Our framework allows users to easily replace the integration scheme to suit their requirements. It is worth noting that in contact-rich simulations involving soft robotic structures, the implicit solver may sometimes struggle to converge. To improve convergence in such scenarios, we provide an optional line search strategy that determines an appropriate step size, ensuring the error decreases with every iteration.
@@ 7.1.  Simulation Experiments
Figure 4: PneuNet actuator modeled as a discrete elastic rod. (a) Snapshots of the PneuNet actuator with natural curvature $15.7$ m-1 with only gravity (top) and with gravity and point force on the free end (bottom). (b) Snapshots of the PneuNet actuator with natural curvature $31.45$ m-1 with only gravity (top) and with gravity and point force on the free end (bottom). (c) Snapshots of the PneuNet actuator with natural curvature $47.15$ m-1 with only gravity (top) and with gravity and point force on the free end (bottom). Note that the cross-sectional radius of the rod is 11 mm, it has been enlarged in the figure for better visualization. (d) The plot of the natural shapes of the rods in (a), (b), and (c) under gravity. (e) Plot of the deformed shape of the rods under gravity and point force at the free end.
[…]
Figure 5: An earthworm modeled as an elastic rod in MAT-DiSMech. Snapshots of the earthworm’s motion at different time stamps (a1-e1) with friction ($\mu=0.25$), and (a2-e2) without friction. Here, time is given in seconds. (f) Displacement of the $x$-coordinate of the front end of the earthworm with time. Notice how, when there is no friction, the earthworm doesn’t move forward, and the contraction and expansion of the edges are also less pronounced compared to when there is friction. The time stamps corresponding to the snapshots (a-e) are marked in the plot using vertical dashed lines.
@@
Snake. As our fourth example, we simulate the serpentine locomotion of a snake swimming in a fluid environment, subject to hydrodynamic forces. The snake is modeled as an elastic rod with a length of $0.1$ m, a cross-sectional radius of $1$ mm, Young’s modulus $E=2$ MPa, Poisson’s ratio $\nu=0.5$, and density $\rho=1200$ kg/m3. The surrounding fluid exerts hydrodynamic forces computed using resistive force theory (Gray et al. [42]), with tangential and normal drag coefficients set to $0.01$ and $0.1$, respectively. This anisotropy in drag coefficients enables forward propulsion when serpentine undulations are applied. Serpentine motion is generated by modulating the natural curvature, $\bar{\boldsymbol{\kappa}}$, of the rod’s bending-twisting springs to produce a sinusoidal wave along the snake’s body, with an amplitude of $0.05$ m and a frequency of $1$ Hz. Figure 7 presents snapshots of the snake in motion and the trajectory of its leading end. The plots show that the $x$-coordinate steadily increases over time as the snake propels forward, the $y$-coordinate oscillates sinusoidally reflecting the serpentine path, while the $z$-coordinate remains constant at zero.
Figure 7: Snake modeled as an elastic rod in MAT-DiSMech. (a-e) Snapshots of the Snake’s motion at different time stamps. Here the time is given in seconds. (f) Displacement of the leading end of the snake with time. The time stamps corresponding to the snapshots (a-e) are marked in the plot using vertical dashed lines.
Parachute. For our fifth showcase, we simulate a parachute which is modeled as a combination of elastic rods and shell. The canopy is modeled as an initially flat plate in the shape of a regular hexagon of side length $1$ m and thickness $1$ mm. The ropes are modeled as elastic rods. A rigid payload mass of $0.13$ kg is attached at the end of the ropes, which is modeled simply by increasing the mass at the node shared between all rods. The material parameters used in this simulation are: $E_{\textrm{rod}}=10$ MPa, $E_{\textrm{shell}}=1$ GPa, $\nu_{\textrm{rod}}=0.5,\nu_{\textrm{shell}}=0.3$ and $\rho=1500$ kg/m3 for both rod and shell. The external forces, in this case, are gravity and aerodynamic drag. We use $\mathbf{g}=[0,0,-9.8]^{\top}$ m/s2, and, $\rho_{\textrm{med}}=1$ kg/m3 and $C_{D}=10$ to calculate the gravity and drag force respectively. Figure 8 shows the snapshots of the parachute as it falls to the ground. It can be seen that as the simulation progresses, the initially flat canopy takes a 3D inverted bowl-like shape as expected.
@@
Rod dropped on ground. For our sixth example, we simulate a rod falling onto the ground from a height of $5$ cm. The density of the rod’s material and its Poisson ratio are taken to be $1500$ kg/m3 and $0.3$, respectively. We simulate the rod for two different values of Young’s modulus $Y$: (a) $Y=2$ MPa and (b) $Y=2$ GPa. The external forces, in this case, are gravity, ground contact, and friction. The values of $0,0,-9.8]^{\top}$ m/s2, $0.25$, and $20$ are used respectively for the acceleration due to gravity $\mathbf{g}$, and the friction coefficient $\mu$ and contact stiffness $k_{c}$ between the ground and the rod. Figure 9 shows snapshots of the rod as it hits the ground. As Young’s modulus for the second case is much higher than the first case, we observe significant deformation of the rod in the first case while almost no deformation in the second.
Figure 9: Elastic rod with different Young’s Moduli dropped on the ground. (a1-f1) Snapshots of the elastic rod with Young’s modulus $2$ MPa, and (a2-f2) Snapshots of the elastic rod with Young’s modulus 22 GPa. Here time is given in seconds. Note that the cross-sectional radius of the rod is 11 mm, it has been enlarged in the figure for better visualization. (g) The trajectory of the $z$-coordinate of the lowermost point on the rod with time. The time stamps corresponding to the snapshots (a-f) are marked in the plot using vertical dashed lines.
Rod gripping a sphere. For our final example, we demonstrate a rod gripping a fixed spherical object. The rod has material properties of $E=200$ MPa, $\nu=0.5$, and $\rho=1200$ kg/m3. It measures $0.1$ m in length with a cross-sectional radius of $0.001$ m. Initially positioned along the $x$-axis, the rod is clamped at one end at the origin, while the sphere is located at the point $0.1,0.0,0.05]^{\top}$. Actuation is performed by gradually varying the natural curvature of the bending-twisting springs along the final three-quarters of the rod’s length, transitioning from $0,0]^{\top}$ to $0.1886,0]^{\top}$. The external forces in this case arise from contact and friction between the rod and the sphere. The simulations use a contact stiffness of 20 and a coefficient of friction of 0.25. Figure 10 illustrates intermediate configurations of a clamped slender rod undergoing actuation via changes in its natural curvature. In case (a), with contact disabled, the rod adopts its natural shape without regard for the spherical obstacle, passing through it if necessary. In contrast, case (b) demonstrates that when contact forces are included through our penalty energy formulation, the rod successfully wraps around the sphere.
Figure 10: Intermediate configurations of a clamped slender rod actuated by varying its natural curvature to bend inward. The rod’s color transitions from yellow to red to indicate progression over time. (a) Without contact modeling, the rod passes through the spherical object to achieve its natural curvature-driven shape. (b) With contact and friction forces modeled via our penalty energy method, the rod interacts with the sphere and successfully wraps around it.
In addition to these examples, prior studies have demonstrated that the modeling techniques utilized in our simulator can capture significantly more complex phenomena such as flagella buckling [8], knot tying [20], and shear-induced pitchfork bifurcations in ribbons [44]. While our examples in this paper are not intended to precisely replicate specific biological trajectories, they are deliberately chosen to showcase the simulator’s ability to model sophisticated motions and handle challenging interactions with the environment – including contact dynamics, frictional forces, and hydrodynamic resistance.